Builder Products, Inc., uses the weighted average method in its process costing system. It manufactures a caulking compound tRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the cost per equivalent unit for materials and conversion forRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the cost of ending work in process inventory for materials, coRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the cost of units transferred out to the next department for mRequired 1 Required 2 Required 3 Required 4 Required 5 Prepare a cost reconciliation report for May. (Round your intermediate
Answer #1

1) Equivalent unit

Material Conversion Cost
Unit transferred out 340000 340000
Ending WIP 25000*80% = 20000 5000
Equivalent unit 360000 345000

2) Cost per equivalent unit

Material Conversion
Beginning WIP 73200 24400
Cost added 380400 134300
Total Cost 453600 158700
Equivalent unit 360000 345000
Cost per equivalent unit 1.26 0.46

3) Cost of ending WIP

Material Conversion Total
Cost of ending WIP 20000*1.26 = 25200 5000*.46 = 2300 27500

4) Cost of unit transferred out

Material Conversion Total
Cost of unit transferred out 340000*1.26 = 428400 156400 584800

5) Cost reconciliation

Cost to be accounted for
Beginning WIP 97600
Cost added 514700
Total Cost to be accounted for 612300
Cost accounted as
Cost of unit transferred out 584800
Cost of ending WIP 27500
Total Cost accounted as 612300